OpenStudy (anonymous):

A card is drawn from a deck of 52. Find the probability of drawing a red card (hearts or diamonds) or a face card (jack, queen, or king).

OpenStudy (anonymous):

I thought that is was 6/13 but that wasn't it. If you could please help i would appreciate it.

OpenStudy (kirbykirby):

Since half the deck is red, the probability of getting a red card = 1/2 There are 4 faces cards of each suit (4 Jacks, 4 Queens, 4 Kings), so the probability of picking one of those is = (4+4+4)/52 = 12/52 So, the probability of getting a red card OR face card means you must use the "addition" rule, so add the 2 probabilities: 1/2 + 12/52 = 13/26 + 6/26 = 19/26